[ ]/18 = 2/9 [ ]/18 = 2/9 [ ]/18 = 2/9 [ ]/18 = 2/9

 [ ]/18 = 2/9
 [ ]/18 = 2/9
 [ ]/18 = 2/9
 [ ]/18 = 2/9

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the missing number be y.
Then,
y/18 = 2/9
By cross multiplication we get,
y = (2 × 18)/9
y = 36/9
Divide both numerator and denominator by 9.
y =4
Therefore,
[4]/18 = 2/9

Thanks